# Uncertainty-Aware Data-Efficient AI: An Information-Theoretic Perspective

### v1

@@ IV-A Set Prediction
Set-valued predictions provide a convenient and actionable way to quantify uncertainty in supervised learning settings. Specifically, given an input $x$, a set-valued predictor $\mathcal{C}(x)$ maps the input to a subset of the output space. For example, in Fig. 1(c), the set $\mathcal{C}(x)$ represents the confidence region (error bars) around the trained predictor, shown as the red shaded area. For a desired miscoverage rate $\alpha\in(0,1)$, e.g., $10\%$ or $\alpha=0.1$, a useful property for the set predictor is to satisfy the coverage guarantee
@@
Moreover, considering any fixed input $x$, one expects that regions of the input space with a higher-entropy distribution $p(y|x)$ yield larger prediction sets. Writing, with some abuse of notation with respect to Table I, as $\mathrm{H}(y|x)$ the entropy of the random variable $y\sim p(y|x)$ for a fixed $x$, this observation can be formalized as follows. Given two inputs $x$ and $x^{\prime}$, if the prediction set $C^{*}(x)$ is always larger than the corresponding set $C^{*}(x^{\prime})$ for all miscoverage levels $\alpha$, the probability vector with entries $\{p(y|x)\}_{y}$ is majorized by the vector $\{p(y|x^{\prime})\}_{y}$ [37]. This, in turn, implies the inequality $\mathrm{H}(y|x)\geq\mathrm{H}(y|x^{\prime})$ between the entropies of the random variables $y\sim p(y|x)$ and $y^{\prime}\sim p(y|x^{\prime})$ [37, 8].
@@ IV-B Conformal Prediction
The most computationally efficient variant is split CP. The procedure partitions the available data into training set $\mathcal{D}$ and calibration set $\mathcal{D}_{\text{cal}}=\{(x_{i},y_{i})\}_{i=1}^{m}$, and it applies the following steps:
Step 1: Training. Train a model $q(y|x)$ using training set $\mathcal{D}$. Note that the model may be trained using conventional frequentist learning or via Bayesian methods, producing an ensemble predictor as in (3).
@@
where ${\mathrm{TV}}(\cdot,\cdot)$ is the total variation distance between two distributions (see Table I). Accordingly, a distribution shift, captured by the total variation distance ${\mathrm{TV}}(p(z_{i}),p(z))$, causes an increase in the miscoverage rate beyond the target value $\alpha$ (see Fig. 2). This issue can be mitigated by assigning calibration data points $z_{i}$ with large expected data shifts a smaller weight $w_{i}$. For example, for a time series, it may be useful to give larger weights to more recent samples.
@@ IV-D On the Efficiency of Prediction Sets
where $m$ is the calibration set size, $\alpha$ is the target miscoverage rate, and $\text{gen}(\theta,\mathcal{D})$ is the generalization error of the pre-trained predictor used to evaluate the scores (see Sec. III). The function $f$ is increasing in the generalization error $\text{gen}(\theta,\mathcal{D})$ and decreasing in both $m$ and $\alpha$.
@@ V-A Prediction-Powered Inference
As illustrated in Fig. 3, prediction-powered inference (PPI) [17, 60, 61] assumes the availability of a small labeled dataset $\mathcal{D}=\{z_{i}=(x_{i},y_{i})\}_{i=1}^{n}$ with true outcomes $y_{i}$, and a much larger dataset $\mathcal{D}_{\mathrm{unl}}=\{(\tilde{x}_{j})\}_{j=1}^{N}$ with inputs $\tilde{x}_{j}$ following the same distribution as the labeled dataset, where $N\gg n$. The goal is to estimate a vector $\theta$, such as the parameters of an AI model [18, 62], using both datasets to achieve better performance than using labeled data alone.
@@ V-B Synthetic-Powered Predictive Inference
As seen in Fig. 3, while PPI leverages synthetic labels to enhance parameter estimation, synthetic-powered predictive inference (SPI) [19], and its generalized form general synthetic-powered predictive inference (GESPI) [20], integrate large synthetic data $\mathcal{D}_{\textup{synth}}$ into CP to improve the informativeness of prediction sets when the real, gold standard calibration data $\mathcal{D}_{\textup{cal}}$ is scarce. The synthetic data can originate from various sources, including unlabeled data with pseudo-labels, generative AI models, or auxiliary tasks.

### v2

@@ III-C Bounding the Generalization Error with High Probability
PAC-Bayes bounds have emerged as one of the most effective theoretical tools for studying generalization in modern deep learning. For example, references [26, 47] demonstrated non-vacuous generalization bounds for deep neural networks and LLMs, respectively. These results support the broader thesis, articulated by [75], that phenomena such as benign overfitting, double descent, can be rigorously characterized by long-standing frameworks like PAC-Bayes that reward flexible hypothesis spaces endowed with soft inductive biases favoring simpler, more compressible solutions.
@@ IV-A Set Prediction
Set-valued predictions provide a convenient and actionable way to quantify uncertainty in supervised learning settings. Specifically, given an input $x$, a set-valued predictor $\mathcal{C}(x)$ maps the input to a subset of the output space. For example, in Figure 1(c), the set $\mathcal{C}(x)$ represents the confidence region (error bars) around the trained predictor, shown as the red shaded area. For a desired miscoverage rate $\alpha\in(0,1)$, e.g., $10\%$ or $\alpha=0.1$, a useful property for the set predictor is to satisfy the coverage guarantee
@@
Moreover, considering any fixed input $x$, one expects that regions of the input space with a higher-entropy distribution $p(y|x)$ yield larger prediction sets. Writing, with some abuse of notation with respect to Table I, as $\mathrm{H}(y|x)$ the entropy of the random variable $y\sim p(y|x)$ for a fixed $x$, this observation can be formalized as follows. Given two inputs $x$ and $x^{\prime}$, if the prediction set $C^{*}(x)$ is always larger than the corresponding set $C^{*}(x^{\prime})$ for all miscoverage levels $\alpha$, the probability vector with entries $\{p(y|x)\}_{y}$ is majorized by the vector $\{p(y|x^{\prime})\}_{y}$ [49]. This, in turn, implies the inequality $\mathrm{H}(y|x)\geq\mathrm{H}(y|x^{\prime})$ between the entropies of the random variables $y\sim p(y|x)$ and $y^{\prime}\sim p(y|x^{\prime})$ [49, 67, 73].
@@ IV-B Conformal Prediction
The most computationally efficient variant is split CP [4]. The procedure partitions the available data into training set $\mathcal{D}$ and calibration dataset $\mathcal{D}_{\text{cal}}=\{(x_{i},y_{i})\}_{i=1}^{m}$, and it applies the following steps:
Step 1: Training. Train a model $q(y|x)$ using training dataset $\mathcal{D}$. Note that the model may be trained using conventional frequentist learning or via Bayesian methods, producing an ensemble predictor as in (3).
@@
where ${\mathrm{TV}}(\cdot,\cdot)$ is the total variation distance between two distributions (see Table I). Accordingly, a distribution shift, captured by the total variation distance ${\mathrm{TV}}(p(z_{i}),p(z))$, causes an increase in the miscoverage rate beyond the target value $\alpha$ (see Figure 2). This issue can be mitigated by assigning calibration data points $z_{i}$ with large expected data shifts a smaller weight $w_{i}$. For example, for a time series, it may be useful to give larger weights to more recent samples.
@@ IV-D On the Efficiency of Prediction Sets
where $m$ is the calibration dataset size, $\alpha$ is the target miscoverage rate, and $\text{gen}(\theta,\mathcal{D})$ is the generalization error of the pre-trained predictor used to evaluate the scores (see Section III). The function $f$ is increasing in the generalization error $\text{gen}(\theta,\mathcal{D})$ and decreasing in both $m$ and $\alpha$.
@@ V-A Prediction-Powered Inference
As illustrated in Figure 3, prediction-powered inference (PPI) [5, 8, 84] assumes the availability of a small labeled dataset $\mathcal{D}=\{z_{i}=(x_{i},y_{i})\}_{i=1}^{n}$ with true outcomes $y_{i}$, and a much larger dataset $\mathcal{D}_{\mathrm{unl}}=\{(\tilde{x}_{j})\}_{j=1}^{N}$ with inputs $\tilde{x}_{j}$ following the same distribution as the labeled dataset, where $N\gg n$. The goal is to estimate a vector $\theta$, such as the parameters of an AI model [83, 65], using both datasets to achieve better performance than using labeled data alone.
@@
The framework of reference [24] provides a useful lens through which to assess the role of PPI in leveraging synthetic data in the context of scaling laws for AI. Their analysis shows that training on a mixture of real and synthetic data introduces an extra error term that scales with the distributional mismatch with real data. This term creates an irreducible floor in the test error that persists unless essentially all synthetic data is discarded. Even reweighting schemes cannot overcome this barrier, as the optimal weight assigned to synthetic data vanishes in the scaling-laws regime. PPI offers a structurally different approach: rather than treating synthetic and real labels as interchangeable, PPI uses a small real dataset to estimate and correct the systematic bias of a synthetic which to assess the role of PPI in leveraging synthetic data. The same principle underlies synthetic-powered predictive inference discussed in the next subsection.
@@ V-B Synthetic-Powered Predictive Inference
As seen in Figure 3, while PPI leverages synthetic labels to enhance parameter estimation, synthetic-powered predictive inference (SPI) [13], and its generalized form general synthetic-powered predictive inference (GESPI) [12], integrate large synthetic data $\mathcal{D}_{\textup{synth}}$ into CP to improve the informativeness of prediction sets when the real, gold standard calibration data $\mathcal{D}_{\textup{cal}}$ is scarce. The synthetic data can originate from various sources, including unlabeled data with pseudo-labels, generative AI models, or auxiliary tasks.